an SRS of 75 incoming calls to a company's customer-service agents showed it took an average of 42.2 seconds for the agents to answer the calls with a standard deviation of 13.20 seconds. Give a 95 percent interval estimate for the mean amount of time it takes the agents to answer calls
Answer by stanbon(75887)

an SRS of 75 incoming calls to a company's customer-service agents showed it took an average of 42.2 seconds for the agents to answer the calls with a standard deviation of 13.20 seconds. Give a 95 percent interval estimate for the mean amount of time it takes the agents to answer calls
-----------
sample mean = x-bar = 42.2 sec
-------
Margin of Error = ER = 1.96*13.2/sqrt(75) = 2.987
------
95% CI: 42.2 - 2.987 < u < 42.2 + 2.987
==============================================
Cheers,
Stan H.
-------------------------